Can you please help me with this? I have to write this equation in slope-intercept form: y+12=-2(x+5/3)
Answer by mananth(16946)

y+12=-2(x+5/3)
y+12= -2(3x+5)/3
multiply by 3
3(y+12)=-6x-10
3y+36=-6x-10
3y=-6x-36-10
3y=-6x-46
/3
y=-2x-(46/3)